OpenStudy (anonymous):

What is the slope of the line that passes through the points (3, –1) and (–2, –5)?

OpenStudy (abb0t):

\(\sf \color{red}{slope = \frac{\Delta x}{\Delta y}}\) = \(\large \sf \color{}{\frac{y_2-y_1}{x_2-x_1}}\)

OpenStudy (anonymous):

A. 07_15_UT_5d.gif B. 07_15_UT_5b.gif C. 07_15_UT_5a.gif D. 07_15_UT_5c.gif

OpenStudy (anonymous):

(3 pt) Select the equation of the line in slope-intercept form: 4x – 9y = –36

OpenStudy (abb0t):

Well, you need to solve the slope first, which is the equation given above. You take the difference between the two points :)